With demand function P = 120 -2Q and MC= 20.
What is the optimal quantity and price for four identical firms.
1
Expert's answer
2017-05-08T08:57:09-0400
P = MC, that`s why 20 = 120-2*Q, Q =50 and P = 20, but If you need find optimal quantity for four firms, then we have that Q = 4*50 = 200.